OpenStudy (anonymous):

working together, a man and a woman can do the job in 8 hours. If either had to do it alone it would take a woman 2 hours longer than a man.How long would it take either person working alone?

OpenStudy (blockcolder):

Let M and W be the # of hours it takes for the man and woman to the job. If they work together, we have \[1/M+1/W=1/8\]. However, if they work alone, we have W=M+2. Plug this in the first equation and solve for M, then W. :D

OpenStudy (anonymous):

\[\left( 1\right)/\left( x \right)+\left( 1 \right)/\left( x+2 \right)=\left( 1 \right)/\left( 8 \right)\]

OpenStudy (anonymous):

how do I solve this?

OpenStudy (blockcolder):

In this case, the solution would go as follows: \[1/M + 1/(M+2)= 1/8\] \[8[(M+2)+M]=M(M+2)\] \[8M+16+8M =M^2+2M\] \[M^2-14M-16=0\] and I don't know what went wrong but M became irrational.

OpenStudy (anonymous):

quadratic equation

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

x= 15.06